# A Simple LLM Framework for Long-Range Video Question-Answering

### v1

@@ 4.2.3 LLM Prompt Analysis
Standard 51.8
Zero-shot Chain-of-Thought [64] 53.2
Plan-and-Solve [57] 54.2
Self-Consistency [59] 55.4
Ours 57.6
Few-shot
Standard 56.5
Ours 61.7
Table 5: Comparison with commonly used prompting techniques. The “Standard” means a standard LVQA prompt is used (see Section 3). We demonstrate that our framework benefits from more sophisticated prompting techniques. Our multi-round summarization-based prompt achieves the best performance in both zero-shot and few-shot learning settings.
Comparison with Commonly Used Prompts. Next, in Table 5, we compare our multi-round summarization-based prompt with other commonly used prompts such as Zero-shot Chain-of-Thought [64], Plan-and-Solve [57], and Self-Consistency [59]. From these results, we observe that all of these prompts outperform the base variant of our model that uses a standard prompt. In particular, among these commonly used prompts, the self-consistency prompting technique achieves the best results (55.4%). Nevertheless, our multi-round summarization-based prompt achieves the best performance (57.6%).
4.2.4 Few-shot In-Context Learning
In-context learning with LLMs has shown strong few-shot performance in many NLP tasks [2, 64]. In Table 5, we evaluate the few-shot in-context learning capabilities of our LLoVi framework. Our results show that our LLoVi framework greatly benefits from few-shot in-context learning. Specifically, the few-shot in-context learning leads to a 4.7% boost on the variant of our framework that uses a standard prompt and 4.1% boost on our advanced framework using a multi-round summarization-based prompt. We used 6 few-shot examples for our experiments as we found this configuration to produce the best performance.
4.3 Main Results on EgoSchema
In Table 6, we evaluate our best-performing LLoVi framework, developed using our empirical insights on the full EgoSchema test set containing 5K video samples. We compare our approach with prior state-of-the-art methods including InternVideo [61], mPLUG-Owl [77], VIOLET [11], FrozenBiLM [73]. Based on these results, we observe that the best-performing zero-shot variant of our LLoVi framework achieves 50.3% accuracy, outperforming the previous best-performing InternVideo model by a large margin (18.2%). Additionally, we show that using few-shot in-context learning, our best variant gets further improvement. These results validate our design choice of leveraging the long-range modeling capabilities of LLMs for the LVQA task. Furthermore, since our proposed LLoVi framework is agnostic to the visual captioning model and an LLM it uses, we believe that in the future, we could further improve these results by leveraging more powerful visual captioners and LLMs.
Model Acc. (%)
@@ 4.3 Main Results on EgoSchema
VIOLET [11] 19.9
FrozenBiLM [73] 26.9
mPLUG-Owl [77] 31.1
InternVideo [61] 32.1
LLoVi (Ours) 50.3
Few-shot
LLoVi (Ours) 52.5
Table 6: Results on the full set of EgoSchema. The best-performing zero-shot variant of our LLoVi framework achieves 50.3% accuracy, outperforming the previous best-performing InternVideo model by 18.2%. For fair comparisons, we gray out our best few-shot variant.
4.4 Results on Other Datasets
 Model Causal (%) Temporal (%) Descriptive (%) All (%)
VFC  [39] 45.4 51.6 64.1 51.5
InternVideo  [61] 43.4 48.0 65.1 49.1
ViperGPT  [51] - - - 60.0
SeViLA  [79] 61.3 61.5 75.6 63.6
LLoVi (ours) 69.5 61.0 75.6 67.7
Table 7: Zero-shot results on NeXT-QA. LLoVi achieves 67.7% accuracy, outperforming previous best-performing model SeViLA by 4.1%. Notably, LLoVi excels at causal reasoning outperforming SeViLA by 8.2% in the causal question category.
Next, we demonstrate that our simple framework generalizes well to other LVQA benchmarks.
NExT-QA. In Table 7, we evaluate LLoVi on the NExT-QA [68] validation set in a zero-shot setting. We compare our approach with prior methods: VFC [39], InternVideo [61], ViperGPT [51], and SeViLA [79]. We observe that LLoVi outperforms the previous best-performing method, SeViLA by 4.1%. Notably, in the Causal category, LLoVi achieves 8.2% improvement. We conjecture this improvement comes from the simple 2-stage design of our LLoVi framework: captioning followed by LLM reasoning. By captioning the video, we are able to directly leverage the reasoning ability of the powerful LLMs and thus achieve good causal reasoning performance.
IntentQA. In Table 8, we evaluate our method on the IntentQA [30] test set. In our comparisons, we include several supervised methods (HQGA [69], VGT [70], BlindGPT [42], CaVIR [29]) and one recent zero-shot approach, SeViLA. From the results in Table 8, we observe that our method greatly outperforms all prior approaches, both in the fully supervised and zero-shot settings.
Model Acc. (%)
Supervised
HQGA [69] 47.7
VGT [70] 51.3
BlindGPT [42] 51.6
CaVIR [29] 57.6
Zero-shot
SeViLA [79] 60.9
LLoVi (ours) 64.0
Table 8: Results on IntentQA. Our zero-shot framework outperforms previous supervised methods by a large margin (6.4%). LLoVi also outperforms the recent zero-shot method, SeViLA, by 3.1%.
NExT-GQA. In Table 9, we extend our framework to the grounded LVQA task and evaluate it on the NExT-GQA [71] test set. We compare LLoVi with the weakly-supervised methods: IGV [32], Temp[CLIP](NG+) [71], FrozenBiLM(NG+) [71] and SeViLA [79]. These baselines are first trained on NExT-GQA to maximize the QA accuracy, and then use ad-hoc methods [71] to estimate a relevant video segment for question-answering. Although LLoVi is not trained on NExT-GQA, it still outperforms these weakly-supervised methods by a large margin on all evaluation metrics. These results demonstrate that in addition to LVQA, our framework can also be used to temporally ground its predictions for more explainable long-range reasoning.
Model mIoP IoP@0.5 mIoU IoU@0.5 Acc@GQA
Weakly-Supervised
IGV [32] 21.4 18.9 14.0 9.6 10.2
Temp[CLIP](NG+) [71] 25.7 25.5 12.1 8.9 16.0
FrozenBiLM (NG+) [71] 24.2 23.7 9.6 6.1 17.5
SeViLA [79] 29.5 22.9 21.7 13.8 16.6
Zero-shot
LLoVi (ours) 37.3 36.9 20.0 15.3 24.3
@@ 4.4 Results on Other Datasets
Table 9: Grounded LVQA results on NExT-GQA. We extend LLoVi to the grounded LVQA task and show that it outperforms prior weakly-supervised approaches on all evaluation metrics. For a fair comparison, we de-emphasize the models that were pretrained on video-language grounding datasets.
@@ 5 Conclusion
In this work, we present a simple, yet highly effective LLM-based framework for long-range video question-answering (LVQA). We thoroughly evaluate various design choices of our approach and use our empirical insights to develop a method that improves upon prior LVQA approaches by a significant margin on the newly introduced EgoSchema benchmark. We also demonstrate that our framework generalizes to other LVQA benchmarks such as NeXT-QA, IntentQA, and it can be extended to grounded LVQA tasks. While our paper does not provide any major technical contributions, we hope that our simple LVQA framework will help inspire new ideas and simplify model design in long-range video understanding.
Appendix A Additional Analysis
A.1 Video Sampling Configurations
In Figure 5, we investigate the sensitivity of LVQA performance on EgoSchem with different video sampling configurations. Specifically, in Subfigure 4(a), we experiment with 4 different clip lengths: 0.5s, 1s, 4s, and 8s. For each clip length, we use the stride that would be sufficient to cover the entire long video input. For these experiments, we use a LaViLa visual captioner and a GPT-3.5 LLM. Our results indicate that LVQA performance is the best when the sampled clip length is 1s. We observe that using an even shorter video clip length (i.e., 0.5s) produces many repetitive/redundant captions, which leads to 2% drop in LVQA performance. Furthermore, we also note that increasing the clip length to longer durations (e.g., 2s-8s) makes the accuracy lower since the extracted captions start to lack detailed visual information needed to answer the question. In addition, in Subfigure 4(b), we fix the clip length to 1s and experiment with 4 different stride values: 1s, 2s, 4s, and 8s. Note that the 1s clips sampled using a 1s stride will cover the entire video without overlap. Our results suggest a gradual decrease in LVQA accuracy when increasing the stride from 1s to 8s. This indicates that having gaps in long video coverage leads to suboptimal LVQA performance. Thus, for the rest experiments, we use 1s video clips sampled with a 1s stride.
((a))
((b))
Figure 5: The Analysis of Video Clip Sampling Strategy on EgoSchema. These results show that sampling 1s video clips with a 1s stride leads to the best LVQA performance.
A.2 Accuracy on Different Question Types
To better understand the strengths and limitations of our LVQA framework, we manually categorize questions in the EgoSchema Subset into 5 categories: (1) Purpose/Goal Identification, (2) Tools and Materials Usage, (3) Key Action/Moment Detection, (4) Action Sequence Analysis, (5) Character Interaction (see Supplementary Materials for details). Note that some questions belong to more than one category. Based on this analysis, we observe that almost half of the questions relate to purpose/goal identification, which makes intuitive sense as inferring human goals/intent typically requires a very long video analysis. We also observe that a significant portion of the questions relate to tool usage, key action detection, and action sequence analysis. Lastly, the smallest fraction of the questions belong to character interaction analysis.
Question Category Category Percentage(%) Acc.(%)
Purpose/Goal Identification 49.2 54.9
Tools and Materials Usage 21.8 50.5
Key Action/Moment Detection 21.6 43.5
Action Sequence Analysis 18.2 52.7
Character Interaction 9.4 63.8
In Table  10, we break down our system’s performance according to each of the above-discussed question categories. Our results indicate that our system performs the best in the Character Interaction category (63.8%). One possible explanation is that the LaViLa model, which we use as our visual captioner, is explicitly pretrained to differentiate the camera wearer from other people, making it well-suited for understanding various interactions between characters in the video. We also observe that our framework performs much worse in the Key Action/Moment Detection category (43.5%). We conjecture that this might be caused by the limitations in the visual captioning model, i.e., if the key action fails to appear in any of the visual captions, the question will be almost impossible to answer. Lastly, we note that our model performs quite well on the remaining categories ($>$50%). It is especially encouraging to see strong results (54.9%) in the Purpose/Goal Identification category since inferring human intentions/goals from the video inherently requires very long-form video analysis.
Appendix B Additional Implementation Details
B.1 Captioners
For most experiments on EgoSchema, we use LaViLa as the visual captioner. For other pre-trained visual captioners, we use off-the-shelf pre-trained models, e.g., BLIP2 [28], EgoVLIP [43].
Dual-encoder. We use TimeSformer [1] base model as the visual encoder and a 12-layer Transformer as the text encoder. The input to the visual encoder comprises 4 RGB frames of size $224\times 224$. We randomly sample 4 frames from the input video clip and use RandomResizedCrop for data augmentation. The video-language model follows a dual-encoder architecture as CLIP [45] and is trained contrastively. Following LaViLa [83], we use 1024 as batch size. We train at a $3\times 10^{-5}$ learning rate for 5 epochs on 32 NVIDIA RTX 3090 GPUs.
Narrator is a visually conditioned autoregressive Language Model. It consists of a visual encoder, a resampler module, and a text encoder. We use the visual encoder (TimeSformer [1] base model) from the pretrained dual-encoder (See the previous paragraph). The resampler module takes as input a variable number of video features from the visual encoder and produces a fixed number of visual tokens (i.e. 256). The text decoder is the pretrained GPT-2 [44] base model with a cross-attention layer inserted in each transformer block which attends to the visual tokens of the resampler module. We freeze the visual encoder and the text decoder, while only training the cross-attention layers of the decoder and the resampler module. Following the design in LaViLa [83], we use a batch size of 256 and a learning rate of $3\times 10^{-5}$. We use AdamW optimizer [22] with $(\beta_{1},\beta_{2})=(0.9,0.999)$ and weight decay 0.01. We train the model on 8 NVIDIA RTX 3090 GPUs for 5 epochs.
Narrating video clips. We use nucleus sampling [16] with $p=0.95$ and return $K=5$ candidate outputs. Then we take the narration with the largest confidence score as the final caption of the video clip.
For NExT-QA, IntentQA and NExT-GQA datasets, we use LLaVA1.5 as the visual captioner and GPT-4 as the LLM. Specifically, we use the llava-1.5-7b-hf variant with the prompt “USER: $<$image$>$. Describe the image in 30 words. ASSISTANT: ”.
B.2 LLMs
For most experiments on EgoSchema we use GPT-3.5 as the LLM. Specifically, we use the gpt-3.5-turbo-0613 variant which has 4K context. When the context length is not enough, we use the gpt-3.5-turbo-16k variant. We use 0 as temperature for all experiments.
We use Llama-2-7b-chat-hf, Llama-2-13b- chat-hf, and Llama-2-70b-chat-hf variants as Llama2 models. For all Llama2 models, we use greedy sampling to generate the output.
B.3 Prompting Techniques Implementation
Prompt Details. We provide detailed prompts for our standard prompt in Table 11, multi-round summarization-based prompt in Table 12, Zero-shot Chain of Thought in Table 13, and Plan-and-Solve prompting in Table 14. To implement Self-Consistency, we set the temperature of GPT3.5 to $0.7$ and run Zero-shot Chain of Thought for 5 times following the design in Self-Consistency [59]. Each run of the model provides a result, and the final output is determined by a majority vote. The prompt for the grounded LVQA benchmark is shown in Table 15.
User Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E). You must not provide any other response or explanation. You are given some language descriptions of a first person view video. The video is 3 minute long. Each sentence describes a clip_length clip. Here are the descriptions: Captions You are going to answer a multiple choice question based on the descriptions, and your answer should be a single letter chosen from the choices. Here is the question: Question Here are the choices. A: Option-A. B: Option-B. C: Option-C. D: Option-D. E: Option-E.   Assistant Answer
Table 11: LLoVi with Standard Prompt on EgoSchema.
User You are given some language descriptions of a first person view video. Each video is 3 minute long. Each sentence describes a clip_length clip. Here are the descriptions: Captions Please give me a num_words words summary. When doing summarization, remember that your summary will be used to answer this multiple choice question: Question.   Assistant Summary
User Please provide a single-letter answer (A, B, C, D, E) to the following multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E). You must not provide any other response or explanation. You are given some language descriptions of a first person view video. The video is 3 minute long. Here are the descriptions: Summary You are going to answer a multiple choice question based on the descriptions, and your answer should be a single letter chosen from the choices. Here is the question: Question Here are the choices. A: Option-A. B: Option-B. C: Option-C. D: Option-D. E: Option-E.   Assistant Answer
Table 12: LLoVi with Multi-round Summarization-based Prompt on EgoSchema. We show the variant (C, Q) normal-$\rightarrow$ S, where we feed the question without potential choices to the summarization stage. Top: caption summarization prompt. Bottom: question answering prompt. In the first stage, GPT3.5 outputs a question-guided summary. In the second stage, GPT3.5 takes the summary without the original captions, then answer the multiple choice question.
User You are given some language descriptions of a first person view video. The video is 3 minute long. Each sentence describes a clip_length clip. Here are the descriptions: Captions You are going to answer a multiple choice question based on the descriptions, and your answer should be a single letter chosen from the choices. Here is the question: Question Here are the choices. A: Option-A. B: Option-B. C: Option-C. D: Option-D. E: Option-E. To answer this question, let’s first prepare relevant information and decompose it into 3 sub-questions. Then, let’s answer the sub-questions one by one. Finally, let’s answer the multiple choice question.   Assistant Sub-questions and Sub-answers   User Please provide a single-letter answer (A, B, C, D, E) to the multiple-choice question, and your answer must be one of the letters (A, B, C, D, or E). You must not provide any other response or explanation. Your response should only contain one letter.   Assistant Answer
Table 14: LLoVi with Plan-and-Solve Prompting on EgoSchema.
Table 15: LLoVi Prompt on NExT-GQA.
@@ B.3 Prompting Techniques Implementation
LaViLa
#C C drops the brick mould.
#O man X moves the cards.
#C C puts the cloth on the table.
#C C moves the dough in the tray.
BLIP2
A person is laying a brick in the dirt.
A child is playing a game of monopoly with a tray of paper plates.
A person is working on a tool.
Woman making dough in a kitchen.
@@
Appendix C Qualitative Analysis
C.1 Captioners
In Table 16 we compare different captions generated by BLIP2 and LaViLa on EgoSchema. LaViLa captions are generally more concise than BLIP2 captions, focusing more on the actions while BLIP2 focuses more on describing the objects. We also observe that LaViLa is better at differentiating the camera wearer and other person. As shown in the second image in Table 16, LaViLa tends to focus more on the action of the other person when the camera wearer and other person both appear in the video.
C.2 LLoVi with Standard Prompt
@@
C.3 LLoVi with Multi-round Summarization-based Prompt

### v3

@@ 4.4 Results on Other Datasets
SeViLA Flan-T5 4B 60.9
LLoVi (ours) GPT-4 1.8T 67.1
Table 9: Results on IntentQA. Our zero-shot framework outperforms previous supervised methods by a large margin (9.5%). LLoVi also outperforms the recent state-of-the-art zero-shot method, SeViLA, by 6.2%.
Method LM Params mIoP IoP mIoU IoU5 Acc
@0.5 @0.5 @ GQA
Weakly-Supervised
IGV - 110M 21.4 18.9 14.0 9.6 10.2
Temp[CLIP] Transformer 130M 25.7 25.5 12.1 8.9 16.0
FrozenBiLM DeBERTa 900M 24.2 23.7 9.6 6.1 17.5
SeViLA Flan-T5 4B 29.5 22.9 21.7 13.8 16.6
@@
LLoVi (ours) GPT-4 1.8T 39.4 38.0 21.5 16.2 26.8
@@
Table 10: Grounded LVQA results on NExT-GQA. We extend LLoVi to the grounded LVQA task and show that it outperforms prior weakly-supervised approaches on all evaluation metrics. For a fair comparison, we de-emphasize the models that were pretrained using video-language grounding annotations.
Next, we demonstrate that our LLoVi framework generalizes well to other LVQA benchmarks.
NExT-QA. In Table 8, we evaluate LLoVi on the NExT-QA Xiao et al. (2021) validation set in a zero-shot setting. We compare our approach with prior methods: VFC Momeni et al. (2023), InternVideo Wang et al. (2022a), ViperGPT Surís et al. (2023), and SeViLA Yu et al. (2023). We observe that the best variant of LLoVi outperforms the previous best-performing method, SeViLA by a significant margin of 10.2%. We conjecture this improvement comes from our decomposition of LVQA into two stages, i.e., short-term captioning followed by long-term reasoning with an LLM, which enables us to harness the power of modern LLMs for this challenging task.
IntentQA. In Table 9, we evaluate our method on the IntentQA Li et al. (2023a) test set. In our comparisons, we include several fully supervised methods (HQGA Xiao et al. (2022a), VGT Xiao et al. (2022b), BlindGPT Ouyang et al. (2022), CaVIR Li et al. (2023b)) and the recent state-of-the-art zero-shot approach, SeViLA. From the results in Table 9, we observe that our method greatly outperforms all prior approaches.
NExT-GQA. In Table 10, we extend our framework to the grounded LVQA task and evaluate it on the NExT-GQA Xiao et al. (2023) test set. To do this, we extract visual captions from each frame and then provide them, along with their corresponding frame indices, to the LLM to identify the required frame indices for answering the question. More details are provided in the Supplementary Material. We compare LLoVi with the weakly-supervised methods: IGV Li et al. (2022), Temp[CLIP](NG+) Xiao et al. (2023), FrozenBiLM (NG+) Xiao et al. (2023) and SeViLA Yu et al. (2023). These baselines are first trained on NExT-GQA to maximize the QA accuracy and then use ad-hoc methods Xiao et al. (2023) to estimate a relevant video segment for question-answering. Although LLoVi is not trained on NExT-GQA, it still outperforms these weakly-supervised methods by a large margin according to all evaluation metrics. These results demonstrate that our framework can be used to temporally ground its predictions for more explainable long-range video understanding.
@@ 5 Conclusion
In this work, we present a simple, yet highly effective LLM-based framework for long-range video question-answering (LVQA). Our framework outperforms all prior models on the newly introduced EgoSchema benchmark. Furthermore, we demonstrate that our approach generalizes to other LVQA benchmarks such as NExT-QA, IntentQA, and it can also be extended to grounded LVQA tasks. Lastly, we thoroughly evaluate various design choices of our approach and analyze the key factors behind the success of our method. We hope that our simple LVQA framework will help inspire new ideas and simplify model design in long-range video understanding.
Limitations
One limitation of our approach is that it might produce suboptimal results if the visual captioning outputs are inaccurate. This might happen because many existing visual captioners suffer from hallucinations and often struggle to effectively capture fine-grained visual details (e.g., fine-grained human-object interactions, etc.). Having said this, our framework is highly flexible and agnostic to the exact visual captioning model that it uses. Thus, we believe that in the future, we will be able to address this limitation by leveraging more powerful visual captioners. Furthermore, another limitation of our approach is that many modern LLMs are not designed for long-context modeling, which is critical for the LVQA task. However, we believe that this limitation will also be addressed in the future via a more sophisticated LLM design, thus, allowing us to incorporate more powerful LLMs for even better LVQA performance.
Acknowledgements
We thank Karttikeya Mangalam, Feng Cheng, Yan-Bo Lin, Yue Yang, and Soumitri Chattopadhyay for their discussion and valuable feedback. This work was supported by the Sony Faculty Innovation Award, Laboratory for Analytic Sciences via NC State University, and ONR Award N00014-23-1-2356.
Appendix A Additional Datasets and Metrics
In this section, we provide detailed information about the datasets and the metrics we use.
  • NExT-QA Xiao et al. (2021) contains 5,440 videos with an average duration of 44s and 48K multi-choice questions and 52K open-ended questions. There are 3 different question types: Temporal, Causal, and Descriptive. Following common practice, we perform zero-shot evaluation on the validation set, which contains 570 videos and 5K multiple-choice questions.
  • IntentQA Li et al. (2023a) contains 4,303 videos and 16K multiple-choice question-answer pairs focused on reasoning about people’s intent in the video. We perform a zero-shot evaluation on the test set containing 2K questions.
  • NExT-GQA Xiao et al. (2023) is an extension of NExT-QA with 10.5K temporal grounding annotations associated with the original QA pairs. The dataset was introduced to study whether the existing LVQA models can temporally localize video segments needed to answer a given question. We evaluate all methods on the test split, which contains 990 videos with 5,553 questions, each accompanied by a temporal grounding label. The metrics we used include: 1) Intersection over Prediction (IoP) Xiao et al. (2023), which measures whether the predicted temporal window lies inside the ground truth temporal segment, 2) temporal Intersection over Union (IoU), and 3) Acc@GQA, which depicts the percentage of accurately answered and grounded predictions. For IoP and IoU, we report the mean values and values with the overlap thresholds of 0.5.
Appendix B Qualitative Analysis
B.1 Visual Captioners
In Table 11, we compare different captions generated by BLIP2 and LaViLa on EgoSchema. LaViLa captions are generally more concise than BLIP2 captions, focusing more on the actions, while BLIP2 focuses more on describing the objects. We also observe that LaViLa is better at differentiating the camera wearer from other people in the video. As shown in the second image in Table 11, LaViLa captions capture the actions of the other people (not just the camera wearer) in the video. In Figure 5, we also visualize 3 EgoSchema videos by displaying 4 sparsely-sampled frames. We observe that our framework using the LaViLa captioner can: 1) differentiate between the camera wearer and other people in the video, 2) assign different character ids to different people, and 3) re-identify people if the video consists of simple interaction between the camera wearer and other people.
@@ B.1 Visual Captioners
LaViLa C drops the brick mould. Man X moves the cards. C puts the cloth on the table. C moves the dough in the tray.
BLIP2 A person is laying a brick in the dirt. A child is playing a game of monopoly with a tray of paper plates. A person is working on a tool. Woman making dough in a kitchen.
@@
((a)) In this 3-minute video, the camera wearer interacts with a man. The camera wearer is always labelled with ’C’ and the man is always labelled as ’X’. Man X appears in the first frame. Even though the video loses track of Man X in the second frame, LaViLa still correctly labels him as ’Man X’ in the last frame.
((b)) The 2-minute video shows multiple people in a shopping mall. LaViLa labels different people with different characters.
((c)) This 30-second video depicts 3 people interacting with each other. Person B appears in the second frame. The thrid frame shows another person interacting with the camera wearer C. Even though person B disappers in the third frame, LaViLa still labels the same entity as Person B in the last frame.
Figure 5: Qualitative captioning results on EgoSchema. Our LaViLa visual captioner can differentiate between the camera wearer and other people by assigning the id ’C’ to the camera wearer and other ids (e.g., ’B’, ’M’, ’X’, ’Y’, etc) to other people. This suggests that our framework using the LaViLa captioner has the basic character ReID ability when the video involves simple interactions between people.
@@
B.2 LLoVi with Standard Prompt
@@
B.3 LLoVi with Multi-round Summarization-based Prompt